OpenStudy (anonymous):

Which is a solution of the equation? 2x + y = 5 A. (2, –1) B. (1, 2) C. (3, –1) D. 5

OpenStudy (anonymous):

to find x intercept sub in 0 wherever u see y

OpenStudy (anonymous):

Just plug in the given values in the given equation and check which will give you 5..

OpenStudy (anonymous):

so 2x+0=5

OpenStudy (anonymous):

@lucyv7129 first read the question properly..

OpenStudy (anonymous):

So it would be 5?

OpenStudy (ipwnbunnies):

No, follow what Water is saying.

OpenStudy (anonymous):

it would be c

OpenStudy (anonymous):

if u sub in the coordinates for c, it gives u 5, which is the answer

OpenStudy (anonymous):

@waterineyes o_0 don't lecture me...

OpenStudy (anonymous):

For example, just put A option in the given equation 2x + y = 5.. Option A is (2, -1) So here x = 2 and y = -1.. Just put these values in 2x + y and check whether it is giving you 5 or not?? for example, first one I am doing it for you.. \[2x + y\] \[2 \color{red}{(2)} + \color{red}{(-1)} \implies 4 - 1 = 3\neq 5\]

OpenStudy (anonymous):

So, Option A is wrong... And just try for other options likewise for B, C and D... And check which will give you 5..

OpenStudy (anonymous):

@lucyv7129 Sorry brother.. You are a scholar how can I deliver Lecture to you...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@waterineyes no need in the capital l in lecture :)

OpenStudy (anonymous):

@waterineyes sure! whatever floats your boat!

OpenStudy (anonymous):

@waterineyes haha, why delete?

OpenStudy (anonymous):

I just saw an option looking like \(\color{red}{Delete}\), so I went for it... :P